What is the activation energy (in J ) of the reaction A B for which the rate constant is 13 sec -1 at 600K and 1.9 sec -1 at 300K
1
Expert's answer
2020-07-29T06:38:44-0400

According to the Arrhenius law, the rate constant and the temperature are related through the activation energy:

"k=A\text{e}^{-\frac{E_a}{k_bT}}".

Therefore, the logarithm of the ratio of the rate constants at temperatures "T_1" and "T_2" is:

"\text{ln}(k_1/k_2) = \frac{E_a}{k_b}(\frac{1}{T_2}-\frac{1}{T_1})".

From this expression , the activation energy will be:

"E_a= k_b\frac{\text{ln}(k_1/k_2)}{(\frac{1}{T_2}-\frac{1}{T_1})}"

"E_a= 1.38\cdot10^{-23}\cdot\frac{\text{ln}(13/1.9)}{(\frac{1}{300}-\frac{1}{600})}=1.6\cdot10^{-20}" J.

Answer: the activation energy is 1.6x10-20 J.


Actually, this energy is in joules per one molecule. If you want the activation energy in J/mol, you must use universal gas constant "R" (8.314 J/( K mol)) instead of Boltzmann constant "k_b" .